Chapter 21, Problem 046 3.60 cm. The charges are 41- 10 19 c what is the value of the net electrostatic force on In the tigure tour partides are fixed along an x axis, separated by distances d particles? 3 ฎ2 e, 3 e, and ฎ4-+8e, with e-1.60 a) particle 1 and b) particle 2 due to the other (a) Number Units (b) Number Units
Answer #1

a) for particle 1

Force on particle 1 by Particle 2 F1= kq1q2 / r^2

= 9*10^9*3*1.6*10^-19*1.6*10^-19 / 0.036^2

= 5.33*10^-25 N ( to the right)

Force on particle 1 by particle 3 F2 = 9*10^9*3*1.6*10^-19*1.6*10^-19 / 0.072^2

= 1.33*10^-25 N ( to the left)

Force on particle 1 by particle 4 F3 = 9*10^9*3*1.6*10^-19*8*1.6*10^-19 / 0.1.08^2

= 4.74*10^-25 N ( to the left)

Net force on Particle 1 = 10^-25 ( 4.74 + 1.33 - 5.33) = 7.4*10^-26 N

b) force on particle 2 by particle 1 = 5.33*10^-25 (to the left)

force on particle 2 by particle 3 = 9*10^9*1.6*10^-19*1.6*10^-19 / 0.036^2

= 1.78*10^-25 N ( To the right)

force on particle 2 by particle 4 = 9*10^9*1.6*10^-19*8*1.6*10^-19 / 0.072^2

= 3.56*10^-25 N ( to the right)

net force on particle 2 = 10^-25(3.56 + 1.78 - 5.33) = 10^-27 N ( to the right)